OpenStudy (liliakarina):

Can anyone give a piece of advice of how determine distance between two points, midpoint coordinates, and coordinates of the point that divides a segment in a given reason?

OpenStudy (anonymous):

m=y2-y1 over x2-x1

OpenStudy (liliakarina):

Is this useful for rectangular coordinates?

OpenStudy (anonymous):

i dont know i think this is for straight lines

OpenStudy (anonymous):

oh its to find the midpoint

OpenStudy (liliakarina):

Yes, I'll post an example

OpenStudy (liliakarina):

If a segment is comprised by the point P1-(2, 0) and P2(8, 5), I need to find the coordinates of te points which divide the segment by the given reasons of r=2/3, r=3/2 and r= -2/3

OpenStudy (anonymous):

yeah im not their yet thats next class lol

OpenStudy (liliakarina):

Don't worry, thanks

OpenStudy (anonymous):

Okay so, there are 2 points. P1=(2,0) P2=(8,5) So let us take, x1=2, y1=0 x2=8, y2=5 We need to find the (x,y) coordinates for which the line segment is divided in ratios of a) 2/3 b) 3/2 c) -2/3 There exists a rule to solve the problem which says, m:n = (mx2 + nx1/m+n, my2+ ny1/m+n) Substituting the values of m=2 and n=3 for (a), we get- 2(8) + 3(2)/2+3 = 16+6/5 = 18/5 and, Substituting the values of m=2 and n=3 for the y-coordinate: 2(5)+3(0)/2+3 = 10+0/5 = 2. So the first set of coordinates which divide the segment in the ratio 2:3 is (18/5, 2). Similarly, for the ratio of 3:2 we apply the formula which says: 3(8) + 2(2)/2+3 = 28/5 and, 3(5) + 3(0)/2+3 = 15/5 = 3 So the coordinates are (28/5, 3) Finally, for the ratio of -2:3 -2(8)+3(2)/3-2 = -16+6/1 = -10 and, -2(5)+3(0)/3-2 = -10/1 = -10 So the coordinates are (-10,-10) which divide the line segment in the ratio of m=-2/3. Basically you just use the Section formula and substitute the values for them, hope this helps!

OpenStudy (liliakarina):

Thanks

OpenStudy (anonymous):

You're welcome! I hope you're clear with the question now :)

OpenStudy (liliakarina):

Of course, :D

OpenStudy (anonymous):

Would you classify mine as a good answer then? xD

OpenStudy (liliakarina):

Totally